OpenStudy (anonymous):

The ages of students in a school are normally distributed with a mean of 15 years and a standard deviation of 2 years. Approximately what percent of the students are between 14 and 18 years old?

OpenStudy (anonymous):

24.17% 62.47% 30.85% 93.32%

OpenStudy (lopus):

\[P(14<x \le18)=P(\frac{ 14-15 }{ 2 }<z \le \frac{ 18-15 }{ 2 })\]

OpenStudy (anonymous):

How do I find the percent of the students from here?

OpenStudy (lopus):

\[P(-0.5<z \le1.5)=P(z \le 1.5)-[1-P(z \le 0.5)]\]

OpenStudy (lopus):

tell me what number you find in the table for 1.5

OpenStudy (anonymous):

0.4332? OK, what do I do now w. this number?

OpenStudy (anonymous):

@lopus ?

OpenStudy (lopus):

this number is z<1.5 later do z<0.5 and replace it in equation

OpenStudy (lopus):

tell me you can solve it